write the slope intercept of the line ( 7,-10) slope -3/4.
Answer by nerdybill(7384)

write the slope intercept of the line ( 7,-10) slope -3/4.
.
The "slope intercept" form of a line is
y = mx + b
where
m is slope
b is the y-intercept
(x,y) are points along that line
.
Your problem gives you:
one point: at ( 7,-10)
m = -3/4
Simply plug in the above and solve for b:
y = mx + b
-10 = (-3/4)(7) + b
-40 = (-3)(7) + 4b
-40 = (-21) + 4b
-19 = 4b
-19/4 = b
.
Our equation is then:
y = (-3/4)x - 19/4 (this is what they're looking for)